OpenStudy (anonymous):

midpoint approximation with 3 subintervals of the integral from 1 to 6 of exp(2sqrt(t))dt

OpenStudy (anonymous):

is that what you are asking \[\int\limits_{1}^{6}e^{\2sqrt{t}}\]

OpenStudy (anonymous):

\[\int\limits_{1}^{6}e ^{2\sqrt{t}}\]

OpenStudy (anonymous):

well...since you want an approximation with 3 intervals...you first have to find the width, since midpoint approximation is cutting the graph into rectangles and adding all the areas

OpenStudy (anonymous):

then you find the midpoint between the intervals. and then multiply the width times the height at each midpoint. You get the answer after you add them all together